OpenStudy (anonymous):

integrate [(x)/(x^2 - 4)]dx

OpenStudy (anonymous):

\[\int\limits_{}^{}(x/x^2 - 4)dx = ?\]

OpenStudy (accessdenied):

Looks like a u-substitution should do the trick: Try: \(u = x^2 - 4\) ?

OpenStudy (anonymous):

thanks, can you finish it please, I can't remember how

OpenStudy (accessdenied):

\( du = 2x\ dx \implies \frac{1}{2}du = x\ dx \) \[ \int \frac{1/2\ du}{u} = \frac{1}{2} \int\frac{du}{u}\] Makes sense?

OpenStudy (anonymous):

you dont see 2 a derivative of x^2 so you need to put 2 up and then divide by 2 as you cant change the original part so it will be 1/2 *ln(x^2 -4) +c

OpenStudy (anonymous):

yes I got it thank you guys so much!

OpenStudy (accessdenied):

You're welcome! :)